A car and a bus started out from the same place, traveling in opposite directions. After 4 hours, they were 384 miles apart. One car was going 8 miles per hour faster than the other. How fast was each car going?
Answer by Paul(988)

Bus's speed = x
Car's speed = x+8
After 4 hours meaning 4 times there speed = distance.
Because distance = time x speed.

4(x+x+8)=384
8x+32=384
8x=352
x=44
44+8=52

Hence the Bus's speed is 44mph, and the car's speed is 52mph.
Paul.